# Solving Modular Linear Systems with a Constraint by parallel decomposition of the Smith form and extended Euclidean division modulo powers of primes divisors

### v2

@@ 1 Introduction
In this paper we consider an extended problem of solving (1) with the constraint that a given linear function $\phi(x)=\langle w,x\rangle$ is coprime to $n$. Such a problem arises in Cryptanalysis, [13, 14], in the problem of inversion of sequences of $m$-tuples over Cartesian products of binary or small prime $p$ fields, $\mathbb{F}_{p}^{m}$ or that of local inversion of non-linear maps over $\mathbb{F}_{p}^{m}$ when the complexity of the sequences is considered in terms of Linear Recurrence Relations (LRR) modulo $p^{m}$. Previous cases of LRR as well as non-linear RRs over the finite field $\mathbb{F}$ are discussed in [12, 13, 14]. For instance when the sequence $\{s_{k},k=0,1,2,\ldots\}$ has minimal linear recurrence, the unknown co-efficients of the minimal polynomial $\{x_{0},\ldots,x_{(m-1)}\}$ satisfy a linear system defined by the Hankel matrix of the sequence an the invertibility of the sequence requires that $x_{0}\neq 0$ which is a special case of a linear function $\ law,x\rangle=x_{0}\neq 0$ over the field of the sequence. When the LRR is considered over modulo $n$ then the corresponding condition becomes $x_{0}$ invertible modulo $n$. Hence the problem considered in this paper is a general problem of solving the system (1) with the constraint that a given linear function $\langle w,x\rangle$ is coprime to $n$.
@@
1.1 Smith form reduction method
Condition for solvability and solutions of the modular system (1) can be obtained in terms of a non-homogeneous integral linear system of the form
@@ 1.1 Smith form reduction method
Find all solutions $x\mbox{ \rm{mod} }n$ when the conditions are satisfied.
We show how this extended problem 1 can be solved by using the Smith form reduction of an augmented matrix, the extended Euclidean division and the CRT.
@@ 1.2 Efficient computation by reduction of Smith form modulo $p^{r}$
What appear to be the main drivers of efficiency, pointed out in this paper, in computation of Smith form for solving the modular system (1) and subsequent checking of existence conditions for solving the constraint on $\phi(x)$ in Problem 1 are
@@
    use of CRT which allows decomposition of the problem into independent sub-problems over the moduli $p^{r(p)}$ corresponding to independent prime divisors of $n$. Hence the problem has a natural parallel computational decomposition.
@@
    further, the Smith form of the resulting augmented matrix is vastly simplified since only GCDs of the form $p^{k}$ are needed. This is because GCD computation of a system of integers $a_{1},a_{2},\ldots,p^{r}$ for a known prime $p$ only involves discovering the largest power $p^{k}$ which divides all the $a_{i}$. Hence explicit Euclidean division by general integers is avoided.
@@
    Another important issue leading to the difficulty of solving modular liner system modulo $n$ arises when $n$ has divisors of the form $p^{r(p)}$ with $p$ prime and $r(p)>1$. For such cases solving the reduced linear systems $Ax=b\mbox{ \rm{mod} }p$ over prime fields $\mathbb{F}_{p}$ does not lead to a solution of the system modulo $n$. Hence in such cases efficient (black box) linear algebra method of [3] cannot be used. Such cases of $r(p)>1$ the modular systems have to be necessarily addressed using the Smith form (or invariant factors) of the matrix $[A,-nI]$. Reduction of Smith form computation over moduli $p^{r}$ for parallel computation has surprisingly not appeared in the past literature.
Due to the above reasons the solution of the problem is greatly simplified by parallel reduction over moduli $p^{r}$ and hence the computations are scalable for large number of unknown variables as well as large size of $n$. In the case when the powers $r(p)=1$ for all prime divisors $p$ of $n$ the reduction to moduli $p$ of the systems lead to independent linear systems over $\mathbb{F}_{p}$. In such a case the individual systems can be solved very efficiently by the Wiedemann’s method known as Black box Linear Algebra [3, 7]. Unfortunately Wiedemann’s method based on minimal polynomial computation of the matrix is not applicable when the matrix is not square and the modulus is $p^{r}$ with $r>1$ (system is not over a field).
@@ 2 Solution of the integral linear system in terms of the Smith form
$\hat{S}\left[\begin{array}[]{ll}S&0\\ 0&0\end{array}\right]$
@@
Denoting $\tilde{x}=Q^{-1}x$ the equation (2) is transformed to the system of equations
@@
in which $x_{2}$ does not appear. Since $S$ is diagonal with invariant factor $f_{i}$ on the $i$-th row, $x_{1}$ is integral iff $f_{i}|\beta_{i}$ for $i=1,2,\ldots,r$. Hence $x_{1}=[d_{1},d_{2},\ldots,d_{r}]^{T}$ and $x_{2}$ an arbitrary $\mathbb{Z}^{l-r}$ tuple gives all the solutions of the transformed equation. Hence all solutions $x$ are
$Q\tilde{x}$
@@ 2.1 Solution of the modular linear system
The modular system (1) has a solution iff there is an integral $\mathbb{Z}^{k}$-tuple $y$ such that
@@
where $y$ is an integral $k$ tuple and $I$ is an $k\times k$ identity matrix. Hence we get an equivalent integral system
@@ 2.2 Solution of the constrained system
Next consider the constraint $\phi(x)=\langle w,x\rangle$ required to be coprime to $n$ for some solution $x$ of (1) or equivalently that of (4). Consider the $(l+k)$ tuple $[w^{T},0^{T}]$ obtained by adding $k$ tuple of zero rows to $w$. Denote
$\tilde{w}=\left[\begin{array}[]{l}w_{0}\\ w_{1}\end{array}\right]=Q^{-T}\left[\begin{array}[]{l}w\\ 0\end{array}\right]$ (7)
@@
$\langle w,x\rangle=\langle\left[\begin{array}[]{l}w\\ 0\end{array}\right],Q^{-1}\left[\begin{array}[]{l}x_{0}\\ x_{1}\end{array}\right]\rangle=\langle w_{0},x_{0}\rangle+\langle w_{1},x_{1}\rangle$
@@ 3.2.1 Solution of the Bezout identity with a prime power dividend
for $x$ modulo $p^{d}$. Once $x$ is solved, $ax=1+p^{y}$ for some $y$ which can be obtained by dividing $ax-1$ by $p^{d}$. Since $a$ is assumed to be in $\mathbb{Z}/(p^{d})$ for any other integer $a$ the Bezout identity solutions $x,y$ can be found by solving $(a\mbox{ \rm{mod} }p^{d})x=1$ modulo $p^{d}$.
@@ 3.4 Using the CRT to solve the modular system
$\begin{array}[]{lcl}P_{j}&=&\prod_{i=1,j\neq i}^{i=m}(p_{i})^{r_{i}}\\ Q_{j}&=&(P_{j})^{-1}\mbox{ \rm{mod} }p_{i}^{r_{i}}\end{array}$
@@ 5 Conclusions
The problem of solving a modular linear system with a given linear function such that the function is coprime to the modulus is solved in terms of the Smith form (giving invariant factors) of an augmented matrix of an integral linear system and evaluation of the linear function modulo prime divisors of the modulus. This procedure is then shown to be efficiently computable in parallel, by decomposing the Smith form computation over prime power divisors of the modulus and operations modulo prime powers. Such a parallel reduction of the modular linear systems modulo prime powers, gives a reduction to the index calculus method of solving the discrete logarithm problem analogous to the Pohlig Hellman reduction for the discrete logarithm computation over cyclic groups when the group order has divisors which are powers of primes. For general modular linear systems the problem with the constraint on the value of a linear function on its solutions is shown to be efficiently solvable by reduction of the smith form computation for integers modulo prime powers and using the CRT to get the solution of the modular system from residues which satisfy the constraint on a linear function modulo prime powers. The arithmetic modulo prime powers is utilized in these computations for solving even the basic identities such as the Bezout identity and is the primary driver of efficiency of computation by parallelization.
@@ Appendix A Byte Arithmetic for Smith form reduction modulo $q^{r}$
To take advantage of efficient byte operations libraries in systems softwares, in the Smith form computation, the solution of the Bezout identity in iterations (12) need to be formulated modulo $q$ where $q=p^{d}$, $d$ is the byte length, instead of modulo $p$. This is explained below for the simplest identity for computing GCD of $(a,q^{r})$ by a unimodular transformation modulo $q=p^{d}$.
Let $p^{g}=\mbox{GCD}(a,q^{r})$. We want to find $x,y$ modulo $q^{r}$ such that $ax+q^{r}y=p^{g}$ by using arithemtic (i.e. addition, multiplictaion and division) modulo $q$. Consider a $q$-adic representation of $a,x,y$ analogous to that in (11)
@@ A.1 Solution modulo $q$
The solution of (18) can now be carried out using the arithmetic modulo $q$ because $(a,q^{r})$ is coprime implies $p^{k}$ does not divide $a$ for any $0<k\leq r$. Hence $a\mod p\neq 0$ as well as $a\mod q\neq 0$. Hence we have unique solutions to the following iterations to solve for the co-efficients of $x^{\prime},y^{\prime}$ in (18).
@@
The iterations (19) require arithemetic modulo $q$ which are byte operations. In a predefined arithmetic library the byte operations are implemented efficiently, for instance by look up tables, for addition and multiplication modulo $q$, inversion modulo $q$. One extra operation required in solving the iterations (19) is the quotient operation by $q$ denoted as $[t_{k}]$ (the carry obtained in evaluating $t_{k}$). Hence modulo $q^{r}$ computations can be carried out efficiently by byte operations which are mod $q$ to solve the Bezout identity and subsequently the Smith form by extending the unimodular matrix computations as shown in Section 4.2.2.

### v3

@@ 1 Introduction
In this paper we consider an extended problem of solving (1) with the constraint that a given linear function $\phi(x)=\langle w,x\rangle$ is coprime to $n$. Such a problem arises in Cryptanalysis, [13, 14], in the problem of inversion of sequences of $m$-tuples over Cartesian products of binary or small prime $p$ fields, $\mathbb{F}_{p}^{m}$ or that of local inversion of non-linear maps over $\mathbb{F}_{p}^{m}$ when the complexity of the sequences is considered in terms of Linear Recurrence Relations (LRR) modulo $p^{m}$. Previous cases of LRR as well as non-linear RRs over the finite field $\mathbb{F}$ are discussed in [12, 13, 14]. For instance when the sequence $\{s_{k},k=0,1,2,\ldots\}$ has minimal linear recurrence, the unknown co-efficients of the minimal polynomial $\{x_{0},\ldots,x_{(m-1)}\}$ satisfy a linear system defined by the Hankel matrix of the sequence an the invertibility of the sequence requires that $x_{0}\neq 0$ which is a special case of a linear function $\langle w,x\rangle=x_{0}\neq 0$ over the field of the sequence. When the LRR is considered over modulo $n$ then the corresponding condition becomes $x_{0}$ invertible modulo $n$. Hence the problem considered in this paper is a general problem of solving the system (1) with the constraint that a given linear function $\langle w,x\rangle$ is coprime to $n$.
@@
1.1 Summery of contribution
Consider a brief summery of contributions made in this paper and the plan of presentation.
  1. 1.
    Solution of the modular system (1) is first addressed in Section 2 using the Smith canonical form for integral matrices. An alternative direct linear algebra method of solving such a system by Gaussian reduction fails even for a simplest system $ax=b\mbox{ \rm{mod} }n$ when $a,n$ are not coprime without first using the Euclidean division to compute the Greatest Common Divisor (GCD) of $a,n$. Solution of the constrained system (3) is obtained in terms of conditions related to the prime divisors of $n$ and the unimodular transformations arising in the Smith form in Theorem 2.
  2. 2.
    The modular system solution can be decomposed as shown in Theorem 4 in terms of its residues by using the CRT by decomposing the system modulo $p^{r(p)}$ for the set of all prime divisors $p$ of $n$. However the resultant systems modulo $p^{r}$ are also not solvable by linear algebra over fields when there is an $r(p)>1$. Such a system then can be solved by the special Smith form for the augmented matrix $[A,-p^{r}I]$ as shown in Theorem 3.
  3. 3.
    Next it is shown in Lemma 3 that the Smith form of $[A,-p^{r}I]$ can be efficiently computed by the solution of the Bezout identity modulo $p^{r}$. This computation involves modulo $p$ divisions only by which the computation of GCD of elements of $A$ and $p^{r}$ is much simpler in computation compared to using Euclidean division to solve the Bezout identity. This is the crux of the contribution of this paper which shows how the direct solution of the Bezout identity modulo $p^{r}$ by $p$-adic arithmetic is more efficient to compute the Smith form for special augmented matrix $[A,-p^{r}I]$ than using the Euclidean division.
  4. 4.
    The above technique leads to the methodology of reducing the Smith form computation of $[A,-nI]$ to $[A,-p^{r}I]$ using only modulo $p$ computation. Hence the methodology is a local prime factor decomposition analogous to the decomposition known as Pohlig-Hellman reduction of the Baby Step Giant Step method of computing the discrete logarithm over cyclic groups. Such a reduction has been missing in the literature since the publication of [2].
  5. 5.
    This paper shows how use of $p$-adic arithmetic for computation of the transform of the augmented matrix $[A,-p^{r}I]$ to Smith form using the Bezout identity results into an efficient computation compared to the Euclidean division and Smith form computation of $[A,-nI]$ by parallel computation of Smith forms of $[A,-p^{r}I]$. Further, the computation of Smith form of $[A,-p^{r}I]$ is also accomplished efficiently for known primes $p$ using $p$-adic arithmetic. This is a computational insight which reverses the usual computation of Bezout identity for the of $\gcd{a,p^{r}}$ by using Euclidean division. It shows that direct $p$-adic arithmetic requires only modulo $p$ computations while Euclidean division to find this GCD requires division with respect to multiple divisors and dividends.
  6. 6.
    In the final section, Section 4, we provide a solution of the constrained linear system (14), (15) over a field.
  7. 7.
    Appendix A, treats the case of solving a linear system modulo $q^{r}$ where $q$ is the power $p^{d}$ of a prime denoting the $p$-adic byte representation of numbers modulo $q^{r}$. The resulting algorithm shows the Smith form computation modulo $q^{r}$ can be carried out in efficient byte operations pre-defined in the computer.
  8. 8.
    For solving modular systems, the finite field linear algebra approach is not applicable when there are prime divisors of $n$ with powers larger than unity. Hence the method proposed in this paper provides a completion of this old problem of linear algebra modulo $n$ by using $p$-adic arithmetic and reducing to moduli $p^{r}$.
Above points briefly summerize contributions of this paper. We focused mainly on the theory begind the computation and have omitted numerical illustrations, which are truly out of scope and would have required a much longer article. This paper is based on the arxiv version [15].
1.2 Linear modular system with a constraint
Condition for solvability and solutions of the modular system (1) can be obtained in terms of a nonhomogeneous Integral Linear System of the form
@@ 1.2 Linear modular system with a constraint
Find a characterization of all solution $x\mbox{ \rm{mod} }n$ when the conditions are satisfied and compute one of the solutions.
In the following sections we show how this extended problem 1 can be solved by using the Smith form reduction of an augmented matrix, the extended Euclidean division and the CRT.
@@ 1.3 Efficient computation by reduction of Smith form modulo $p^{r}$
As briefly pointed out above in Section 1.1, Summery of Contributions, the main drivers of efficiency, in computation of the Smith form for solving the modular system (1) and subsequent checking of existence conditions for solving the constraint on $\phi(x)$ in Problem 1 are
@@
    Use of CRT which allows decomposition of the problem into independent sub-problems over the moduli $p^{r(p)}$ corresponding to coprime divisors of $n$. Hence the solution has a natural parallel computational decomposition.
@@
    Further, the Smith form computation of the resulting augmented matrix is vastly simplified since only gcds of the form $p^{k}$ are needed. This is because gcd computation of a system of integers $a_{1},a_{2},\ldots,p^{r}$ for a known prime $p$ only involves discovering the largest power $p^{k}$ which divides all the $a_{i}$. Hence explicit Euclidean division by general integers is avoided.
@@
    Another important issue leading to the difficulty of solving modular liner system modulo $n$ arises when $n$ has divisors of the form $p^{r(p)}$ with $p$ prime and $r(p)>1$. For such cases solving the reduced linear systems $Ax=b\mbox{ \rm{mod} }p$ over prime fields $\mathbb{F}_{p}$ does not lead to a solution of the system modulo $n$. Hence in such cases, efficient (black box) linear algebra method of [3] cannot be used. Such cases of $r(p)>1$ the modular systems have to be necessarily addressed using the Smith form (or invariant factors) of the matrix $[A,-p^{r(p)}I]$. Reduction of Smith form computation over moduli $p^{r}$ using parallel computation (and CRT) has surprisingly not appeared in the past literature.
Due to the above reasons, the solution of the problem is greatly simplified by parallel reduction over moduli $p^{r}$ and hence the computations are scalable for large number of unknown variables as well as large size of $n$. In the case when the powers $r(p)=1$ for all prime divisors $p$ of $n$ the reduction to moduli $p$ of the systems lead to independent linear systems over $\mathbb{F}_{p}$. In such a case the individual systems can be solved very efficiently by the Wiedemann’s method known as Black box Linear Algebra [3, 7]. Unfortunately, Wiedemann’s method based on minimal polynomial computation of the matrix is not applicable when the matrix is not square and the modulus is $p^{r}$ with $r>1$ (system is not over a field).
@@ 2 Solution of the integral linear system in terms of the Smith form
$\hat{S}=\left[\begin{array}[]{ll}S&0\\ 0&0\end{array}\right]$
@@
Denoting $\tilde{x}=Q^{-1}x$ the equation and transforming the equation (2) to the system of equations $(PAQ)(Q^{-1})x=Pb$ which is
@@
in which $x_{2}$ does not appear. Since $S$ is diagonal with invariant factor $f_{i}$ on the $i$-th row, $x_{1}$ is integral iff $f_{i}|\beta_{i}$ for $i=1,2,\ldots,r$. Hence $x_{1}=[d_{1},d_{2},\ldots,d_{r}]$ and $x_{2}$ an arbitrary $\mathbb{Z}^{l-r}$ tuple gives all the solutions of the transformed equation $\tilde{x}=[x_{1},x_{2}]^{T}$. Hence all solutions are
$x=Q\tilde{x}$
@@ 2.1 Solution of the modular linear system
The modular system (1) has a solution $x$ in $\mathbb{Z}^{l}$ iff there is an integral $k$-tuple $y$ in $\mathbb{Z}^{k}$ such that
@@
where $I$ is an $k\times k$ identity matrix. Hence we get an equivalent integral system
@@ 2.2 Solution of the constrained system
Next consider the constraint $\phi(x)=\langle w,x\rangle$ required to be coprime to $n$ for some solution $x$ of (1) or equivalently that of (4). Consider the $(l+k)$ tuple $[w^{T},0^{T}]^{T}$ obtained by adding $k$ tuple of zero rows to $w$. Denote
$\tilde{w}=\left[\begin{array}[]{l}w_{0}\\ w_{1}\end{array}\right]=Q^{T}\left[\begin{array}[]{l}w\\ 0\end{array}\right]$ (7)
@@
$\langle w,x\rangle=\langle\left[\begin{array}[]{l}w\\ 0\end{array}\right],Q\left[\begin{array}[]{l}x_{0}\\ x_{1}\end{array}\right]\rangle=\langle w_{0},x_{0}\rangle+\langle w_{1},x_{1}\rangle$
@@ 3.2.1 Solution of the Bezout identity with a prime power dividend
for $x$ modulo $p^{d}$. Once $x$ is solved, $ax=1+py$ for some $y$ which can be obtained by dividing $ax-1$ by $p^{d}$. Since $a$ is assumed to be in $\mathbb{Z}/(p^{d})$ for any other integer $a$ the Bezout identity solutions $x,y$ can be found by solving $(a\mbox{ \rm{mod} }p^{d})x=1$ modulo $p^{d}$.
@@
Remark 3.
Another important observation about the computation of the Bezout identity and $Q$ using the iterations (12) is that the successive Euclidean divisions of $a$ by $p^{d}$ are not used. The only equations to be solved at all iteration steps are of the form $a_{0}x=b\mbox{ \rm{mod} }p$. On the other hand Euclidean division method of computing the Bezout identity involves several division by dividends other than $p^{d}$. Hence the direct computation of the Bezout identity using the $p$-adic iterations (12) is far more inexpensive.
@@ 3.4 Using the CRT to solve the modular system
$\begin{array}[]{lcl}P_{j}&=&\prod_{i=1,i\neq j}^{i=m}(p_{i})^{r_{i}}\\ Q_{j}&=&(P_{j})^{-1}\mbox{ \rm{mod} }p_{j}^{r_{j}}\end{array}$
@@ 5 Conclusions
The problem of solving a modular linear system with a given linear function such that the function is coprime to the modulus is solved in terms of the Smith form of an augmented matrix of an integral linear system and evaluation of the linear function modulo prime divisors of the modulus. This procedure is then shown to be efficiently computable in parallel, by decomposing the Smith form computation over prime power divisors of the modulus and operations modulo the primes. Such a parallel reduction of the modular linear systems modulo prime powers, gives a reduction of the index calculus method of solving the discrete logarithm problem analogous to the Pohlig Hellman reduction for the discrete logarithm computation over cyclic groups when the group order has divisors which are powers of primes. For general modular linear systems the problem with the constraint on the value of a linear function on its solutions is shown to be efficiently solvable by reduction of the smith form computation for integers modulo prime powers and using the CRT to get the solution of the modular system from residues which satisfy the constraint on a linear function modulo prime powers. The arithmetic modulo prime powers is utilized in these computations for solving even the basic identities such as the Bezout identity and is the primary driver of efficiency of computation by parallelization.
@@ Appendix A Byte Arithmetic for Smith form reduction modulo $q^{r}$
To take advantage of efficient libraries for byte operations available in systems software, in the Smith form computation, the solution of the Bezout identity in iterations (12) need to be formulated modulo $q$ where $q=p^{d}$, $d$ is the byte length, instead of modulo $p$. This is explained below for the simplest identity for computing GCD of $(a,q^{r})$ by a unimodular transformation modulo $q=p^{d}$.
Let $p^{g}=\mbox{GCD}(a,q^{r})$ where $g\geq d$. We want to find $x,y$ modulo $q^{r}$ such that $ax+q^{r}y=p^{g}$ by using arithmetic (i.e. addition, multiplication and division) modulo $q$. Consider a $q$-adic representation of $a,x,y$ analogous to that in (11)
@@ A.1 Solution modulo $q$
The solution of (18) can now be carried out using the arithmetic modulo $q$ because $(a,q^{r})$ is coprime implies $p^{k}$ does not divide $a$ for any $0<k\leq r$. Hence $a\mbox{ \rm{mod} }p\neq 0$ as well as $a\mbox{ \rm{mod} }q\neq 0$. Hence we have unique solutions to the following iterations to solve for the co-efficients of $x^{\prime},y^{\prime}$ in (18).
@@
The iterations (19) require arithmetic modulo $q$ which are byte operations. In a predefined arithmetic library the byte operations are implemented efficiently, for instance by look up tables, for addition and multiplication modulo $q$, inversion modulo $q$. One extra operation required in solving the iterations (19) is the quotient operation by $q$ denoted as $[t_{k}]$ (the carry obtained in evaluating $t_{k}$). Hence modulo $q^{r}$ computations can be carried out efficiently by byte operations which are mod $q$ to solve the Bezout identity and subsequently the Smith form by extending the unimodular matrix computations as shown in Section 4.2.2.